OpenStudy (anonymous):

solve the rational equation. check your answer. 3/ m-4 + 1/3 (m - 4) = 6/m

OpenStudy (anonymous):

factorise on left hand side and cross multiply

OpenStudy (anonymous):

I don't know how to factor

OpenStudy (zale101):

this my teach u how to factor: If u still need help, tag me http://www.purplemath.com/modules/solvrtnl.htm

OpenStudy (jhannybean):

\[\large \frac{ 3 }{ m-4 }+\frac{ 1 }{ 3(m-4) }=6m\]\[\large \frac{ 3 }{ 3 }*\frac{ 3 }{ (m-4) }+\frac{ 1 }{ 3(m-4) }=6m\] you multiply both 3 on the top and bottom of the first fraction because the LCD = 3(m-4) So you get \[\large \frac{ 3(3) }{ 3(m-4) }+\frac{ 1 }{ 3(m-4) }=6m\]Now you've got your LCD are you able to simplify the rest?

OpenStudy (anonymous):

From the beginning multiply both sides by 3( m-4) in order to eliminate fractions quickly and in one

OpenStudy (anonymous):

No, I don't really understand any of it

OpenStudy (zale101):

What part u didn't get?

OpenStudy (anonymous):

I know my answer needs to be 9, but i don't know howto show my work.

OpenStudy (zale101):

jhanny was on the right track, what part u didn't get from it?

OpenStudy (anonymous):

I understand what she has written so far I just don't know what to do after that.

jhonyy9 (jhonyy9):

so attention on the right side wann being there 6 divide by m and not 6 time m

OpenStudy (jhannybean):

From what I have posted,simplify the expression. \[\large \frac{ 9 }{ 3m-12 }+\frac{ 1 }{ 3m-12 }\] since we have a common denominator, we can simplify further, \[\large \frac{ 10 }{ 3m-12 }=6m\] Now we want all the m's on one side. s we can "cross multiply" \[\large \frac{ 10 }{ 3m-12 }=\frac{ 6m }{ 1}\]\[\large 10=6m(3m-12)\] can you carry on from here?

OpenStudy (anonymous):

I only know how to do the first step. I can try if you tell me what I need to do next.

jhonyy9 (jhonyy9):

@Jhannybean on the right side wann being 6/m and not 6m

OpenStudy (jhannybean):

Thank you @jhonyy9 it was a typo :( Missed that!

jhonyy9 (jhonyy9):

was my pleasure but ATTENTION

OpenStudy (anonymous):

So, what do I need to do next?

jhonyy9 (jhonyy9):

@Jhannybean help for medal please

OpenStudy (jhannybean):

Are you asking for one? haha

OpenStudy (anonymous):

I will give a medal to whoever helps finish the problem

OpenStudy (anonymous):

@jhonyy9 and @Jhannybean

OpenStudy (jhannybean):

so we have \[\large \frac{ 10 }{ 3m-12 }=\frac{ 6 }{ m }\]\[\large 10m=6(3m-12)\]\[\large 10m = 18m-72\]\[\large 72=18m-10m\]\[\large 72=8m\] and now what is \[\frac{ 72 }{ 8 }\]? You will have your answer.

OpenStudy (anonymous):

9, Thank you very much.

OpenStudy (jhannybean):

No problem